OpenStudy (anonymous):

Travis is an avid skier. He is trying to decide whether to buy ski equipment. Typically, it costs him $60 a day to rent ski equipment and buy a lift ticket. You can buy ski equipment for about $400. A lift ticket alone costs $35 for one day. Find the break-even point. a.4 days b.11 days c.6 days d.16 days

OpenStudy (triciaal):

one possible solution let the no of days = d 60*d = 400 + 35*d solve for d d = 16